Find the cosine of the angle between the planes x + y + z = 0 and x + 2y + 5z = 3.

<p>Find the cosine of the angle between the planes x + y + z = 0 and x + 2y + 5z = 3.</p>
Answer #1
we have,
A. B =

1 + 2 + 5 = (3)^0.5 x (30)^0.5 x cos(theta)

angle theta = 32.51 degreees
answered by: larrial
Answer #2

normal vectors to planes

(1,1,1) and (1,2,2)

use formula
VdotU=||V||*||U||*cos(pheta)

VdotU=1*1+1*2+1*2=5
||V||=sqrt(1^2+1^2+1^2=sqrt(3)
||U||=sqrt(1^2+2^2+2^2)=sqrt(9)
cos(pheta)=5sqrt(3)/9
pheta=15.8 degrees

answered by: kehau